OpenStudy (anonymous):

Find the particular solution to the differential equation 6(du)/(dt)=u^2 subject to the initial condition u(0)=8

OpenStudy (anonymous):

well this is an interesting question, but simple

OpenStudy (anonymous):

integrate dx/dt ?

OpenStudy (anonymous):

in order to find x(t)

OpenStudy (anonymous):

then substitute x(t) in y(t)

OpenStudy (anonymous):

hence, differentiate y(t), you get the answer

OpenStudy (anonymous):

alright I'll try it

OpenStudy (anonymous):

this is a bit different than the normal way students are used to in solving parametric equations

OpenStudy (anonymous):

Im a bit confused what I do with dx/dt=3 and x=10

OpenStudy (anonymous):

Hint: when you integrate you'll have a constant, so the number is there to help find that constant

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\int\limits dx = 3 \int\limits dt\]

OpenStudy (anonymous):

\[x = 3t + c\]

OpenStudy (anonymous):

I am sorry but I dont see how you got that

OpenStudy (anonymous):

what level are you at?

OpenStudy (anonymous):

Calculus for business

OpenStudy (anonymous):

well do you agree if you integrate 3 with respect to t you get 3t + c

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well that's all i did x = 3t +c

OpenStudy (anonymous):

ok I see, but now how do I use that toward the equation

OpenStudy (anonymous):

now you plug in x = 3t +c in y

OpenStudy (anonymous):

199 = 3t + c ?

OpenStudy (anonymous):

y = 2(3t+c)^2 - 1

OpenStudy (anonymous):

now find dy/dt

OpenStudy (anonymous):

dy/dt=12t+4c-1?

OpenStudy (anonymous):

sorry no "-1"

OpenStudy (anonymous):

expand the function first

OpenStudy (anonymous):

y=18t^2+4ct+2c^2-1

OpenStudy (anonymous):

I thought you had to find the prime first

OpenStudy (anonymous):

well you can do so but i am trying to make this as simple as possible

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y' = 36t + 12c

OpenStudy (anonymous):

2[9t^2+6ct+c^2]-1

OpenStudy (anonymous):

now the question is how do you find c?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

well i think you could do that, this is done using simple method of linear equation y = mx +c, in your case y = 3t +c, what do you do to find c?

OpenStudy (anonymous):

sorry x = 3t +c

OpenStudy (anonymous):

if t = 0 and x =10 what is c?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

good

OpenStudy (anonymous):

your done!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

your welcome! it is a good question worth practicing and spending time on it

OpenStudy (anonymous):

Yes and I will look over this again after a few more problems, do be sure that I could do it on my own now. Thanks again

OpenStudy (anonymous):

I am pretty sure you can, and that makes me happy all the best